A lender requires a minimum down payment of 12% of the value of the home. You have $19,000 cash available to use as a down payment toward a home. Determine the maximum home value that you can finance.
You can afford to finance a home worth at most

Answer by math_tutor2020(3816)


H = home value in dollars

12% of H = 0.12*H = down payment

$19,000 is the most you can afford as the down payment.
Set this equal to 0.12H and solve for H

0.12H = 19,000
H = 19,000/0.12
H = 158,333.333333333 approximately
H = 158,333.33 dollars is the max home value you can afford.